### v3

# Focus on This, Not That! Steering LLMs With Adaptive Feature Specification

Abstract
Despite the success of Instruction Tuning (IT) in training large language models (LLMs) to perform arbitrary user-specified tasks, these models often still leverage spurious or biased features learned from their training data, leading to undesired behaviours when deploying them in new contexts. In this work, we introduce Focus Instruction Tuning (FIT), which trains LLMs to condition their responses by focusing on specific features whilst ignoring others, leading to different behaviours based on what features are specified. Across several experimental settings, we show that focus-tuned models can be adaptively steered by focusing on different features at inference-time: for instance, robustness can be improved by focusing on core task features and ignoring spurious features, and social bias can be mitigated by ignoring demographic categories. Furthermore, FIT can steer behaviour in new contexts, generalising under distribution shift and to new unseen features at inference time, and thereby facilitating more robust, fair, and controllable LLM applications in real-world environments.
Machine Learning, ICML
1 Introduction
Instruction Tuning (IT) (Zhang et al., 2023), a specialised form of supervised fine-tuning (SFT), has become an essential step in the process of developing effective instruction-following large language models (LLMs) (Ouyang et al., 2022; Touvron et al., 2023; Chen et al., 2024). While extensive pre-training to perform next-token prediction allows LLMs to extract common patterns and knowledge from large text corpora, IT fine-tunes these models on input-output pairs complemented by natural-language task instructions, teaching them to perform open-ended language-based tasks given instructions (Huang et al., 2023).
However, despite the improvements observed in zero-shot generalisation from IT, recent studies suggest that some of these gains may be superficial, stemming from the models’ ability to learn task template formats or spurious input/output correlations rather than a more generalisable instruction-following capability (Kung & Peng, 2023; Ghosh et al., 2024). As a result, LLMs may fail to generalise to new contexts where the same templates or spurious correlations are not present (Kung & Peng, 2023). Furthermore, fine-tuning can inadvertently lead to safety misalignment, where models lose alignment with desired objectives and become more prone to generating harmful or undesirable outputs (Qi et al., 2023). This motivates the need for methods that can adaptively steer models at inference time, enabling users to control model behaviours based on evolving requirements or safety considerations.
Figure 1: Focus Instruction Tuning (FIT). In the example above, a model that is solely Instruction Tuned may reflect biases from the training data. For instance, in a question from BBQ (Parrish et al., 2022), when asked who posed a technical question at an engineering candidate’s interview involving both a man and a woman, the model might incorrectly answer “the man” due to biases, despite evidence to the contrary. In contrast, a FIT model can ignore the gender feature and focus on the interview content, demonstrating steerability and adaptability at inference time.
Accordingly, we propose Focus Instruction Tuning (FIT), an extension of traditional IT that fine-tunes LLMs with respect to an instruction specifying which features to “focus on” or “ignore.” FIT trains LLMs to condition responses based on these focus specifications and respond differently to the same task input based on the specified features, allowing end users to dynamically steer model behaviour. This capability provides precise, explainable control over features leveraged by models, and can be used to enforce desired invariances. For instance, in Figure 1, we illustrate how FIT can be used to steer a model to ignore gender stereotypes and focus on task-relevant information, enabling it to correctly solve a question-answering task.
In our experiments, we demonstrate that FIT enables precise control over model behaviour by steering it to focus on task-relevant features while ignoring features specified as irrelevant or spurious (see Section 4). This flexibility allows FIT to address challenges such as mitigating the influence of demographic features in question-answering tasks, while also proving effective across diverse NLP tasks like sentiment analysis and natural language inference. Furthermore, we show that FIT is robust to distribution shifts over feature values and can generalise to new, unseen features, demonstrating its adaptability to dynamic user requirements and varying task contexts.
In summary, our primary contributions are as follows:
  1. 1.
    We introduce Focus Instruction Tuning (FIT), a method that allows users to flexibly and dynamically specify what features a model should or should not focus on when performing a task at inference time. FIT enables practitioners to incorporate domain-specific knowledge about core, spurious, or bias-relevant features in order to steer models according to the desired feature specification.
  2. 2.
    We experiment with FIT across several key NLP tasks, including sentiment analysis, natural language inference, and question-answering. We find that FIT is highly effective for steering behaviour on all tasks with respect to a variety of lexical, distributional, semantic, and demographic features.
  3. 3.
    We show that focus tuning generalizes both to new features not seen during training, and to distribution shift over feature values.
2 Background and Related Work
2.1 Spurious Feature Learning
Deep neural networks, such as foundation models like LLMs, are susceptible to relying on spurious features present in the training dataset – i.e., input features that are correlated with outputs in the training distribution, but are not correlated in all test distributions (Ye et al., 2024). Relying on spurious features leads models to fail to generalise under distribution shifts where such correlations may no longer hold (Wang et al., 2023a). Spurious features have been extensively studied in computer vision, encompassing features such as background colour (Arjovsky et al., 2019; Xiao et al., 2021; Venkataramani et al., 2024; Hemmat et al., 2024), texture (Geirhos et al., 2018; Baker et al., 2018), or scene elements (Hemmat et al., 2024), and are also prevalent in many widely used NLP benchmarks (Sun et al., 2024; Borkan et al., 2019). For instance, the token SPIELBERG is spuriously correlated with positive sentiment in datasets like SST-2 (Socher et al., 2013b), meaning that models trained on SST-2 may learn to predict sentiment by leveraging these spurious features instead of more general sentiment features (Wang & Culotta, 2020). This reliance on non-causal features undermines the robustness of models in generalising to distribution shift. Traditional approaches for detecting and mitigating spurious feature learning, particularly under distribution shifts, include prompt engineering (Sun et al., 2024), regularisation techniques (Arjovsky et al., 2019; Chew et al., 2024), counterfactual inference (Wang & Culotta, 2020, 2021; Udomcharoenchaikit et al., 2022), or generating synthetic interventional data (Bansal & Grover, 2023; Yuan et al., 2024; Wang et al., 2024).
Mechanistic Interpretability.
Substantial work in mechanistic interpretability has also aimed to discover models’ latent representation of, and reliance on, various features (Davies & Khakzar, 2024). For instance, causal probing trains supervised probing classifiers to predict and modify feature representations encoded by foundation models Belinkov, 2022, and has been deployed to study how LLMs leverage task-causal versus spurious features (Ravfogel et al., 2021; Lasri et al., 2022; Davies et al., 2023; Canby et al., 2024). Other works have leveraged unsupervised mechanistic interpretability methods, such as circuit discovery techniques (Wang et al., 2023b; Conmy et al., 2023) and sparse auto-encoders (Subramanian et al., 2018; Yun et al., 2021), to improve generalisation by discovering spurious features leveraged by models in performing a given task and ablating their use of these features (Gandelsman et al., 2024; Marks et al., 2024). Finally, concept removal methods locate and manipulate supervised feature representations corresponding to bias features encoded by foundation models in order to remove these features (Ravfogel et al., 2020, 2022, 2023; Iskander et al., 2023; Belrose et al., 2024; Kuzmin et al., 2024).
2.2 Controlling LLMs
Instruction Tuning. Due to the next-word prediction training objective, large language models (LLMs) often struggle by default to generate outputs that align with human instructions in downstream applications (Huang et al., 2023). Instruction-tuning (IT) mitigates this issue by fine-tuning pre-trained LLMs on datasets composed of instruction-response pairs (Zhang et al., 2023), aiming to align the responses of the fine-tuned model more closely with the distributions preferred by humans (Ouyang et al., 2022). There are several popular approaches for collecting IT training data, such as using human-annotated data (Dolly, 2023), extracting datasets from existing collections (Longpre et al., 2023; Mishra et al., 2022), or gathering data from internet sources (Zhou et al., 2024). IT datasets can also be synthesised with LLMs, either by bootstraping them from the same model that will be instruction-tuned on them (Wang et al., 2023c; Chen et al., 2024), or by distilling from a larger or more powerful model to instruction-tune smaller models (Taori et al., 2023; Mitra et al., 2023; Xu et al., 2023).
Despite the success of IT in zero-shot generalisation, Gudibande et al. (2023) find that improvements on many downstream benchmark tasks may be largely due to coverage of task data within IT training datasets; and bootstrapping IT methods (which, in principle, might not be subject to this issue provided they synthesise novel IT task instances) require a robust and effective LLM for fine-tuning to avoid degenerate training cycles (Zhang et al., 2023). Furthermore, Kung & Peng (2023) show that some of the downstream performance gains from IT can be attributed to models’ ability to learn surface-level patterns, such as the required answer format, rather than acquiring more generalisable instruction-following skills. These limitations underscore the need for advancements in supervised fine-tuning (SFT) methods beyond IT to facilitate more predictable and reliable control of downstream model behaviours.
Aligning LLMs. Alignment techniques like Reinforcement Learning with Human Feedback (RLHF) (Bai et al., 2022) are powerful tools for aligning LLMs with annotated preference data and lead to reduced prevalence of harmful behaviour (Ouyang et al., 2022; Bai et al., 2022; Touvron et al., 2023; Korbak et al., 2023). However, RLHF-trained models still exhibit key alignment limitations such as sycophancy (defaulting to agreement with users even when incorrect or harmful; Perez et al., 2023; Sharma et al., 2024), and can still be adversarially prompted to generate harmful responses (Carlini et al., 2024). Furthermore, even well-aligned models can rapidly fall out of alignment when they are fine-tuned (Zhan et al., 2024; Yang et al., 2024; Lermen & Rogers-Smith, 2024), even on benign tasks (Qi et al., 2023). Thus, effectively steering model behaviours across the range of potential safety concerns that might emerge during LLM pre-training, fine-tuning, or deployment to novel contexts is an important and challenging goal in AI safety, necessitating more flexible and generalisable steering methods (Anwar et al., 2024).
Latent Steering.
A growing literature has worked to address this challenge via inference-time steering, where LLMs do not need to be retrained with respect to safety limitations, but can instead be controlled at inference time to steer models towards desirable behaviours or away from undesirable ones. For instance, latent steering methods perform embedding-space interventions that push models towards desirable behaviours or away from undesirable ones (Turner et al., 2023; Zou et al., 2023; Bhattacharjee et al., 2024; Li et al., 2024; Han et al., 2024). However, these methods require white-box access to model representations at inference time; and interventions are specific to each target behaviour, meaning that they must be re-computed in the context of any given target behaviour. Our work addresses these limitations by training models to control and adapt their responses based on explicit focus specifications, enabling users to flexibly and dynamically steer model behaviour at inference-time using simple, natural-language instructions.
3 Methodology
Preliminaries. We consider a pre-trained, decoder-only large language model (LLM) $p_{\theta}$ that models the probability of token sequences autoregressively over its vocabulary $\mathcal{V}$. Given a sequence of tokens $s=[s_{1},\dots,s_{L}]\in\mathcal{V}^{L}$, the jonit probability of $s$ under the model is given as
$p_{\theta}(s)\;=\;\prod_{i=1}^{L}\,p_{\theta}(s_{i}\mid s_{<i}),$ (1)
where $p_{\theta}(s_{1}\mid\emptyset)=p_{\theta}(s_{1})$. In supervised fine-tuning, we minimise the negative log-likelihood (NLL) of the output tokens $y$ given the input tokens $x$ using the autoregressive formulation defined in Equation 1.
In instruction tuning (IT) (Zhang et al., 2023), a form of SFT, an additional task instruction $I$ accompanies the input-output pair $(x,y)$, forming a tuple $(I,x,y)$. The objective becomes minimising the NLL of $y$ given both $I$ and $x$ over the distribution of input-output pairs, and instructions.
Focus Instruction Tuning (FIT). We introduce Focus Instruction Tuning (FIT), a specialised form of instruction tuning that trains LLMs to adjust their responses based on user-specified features provided in natural language.
Let $\mathcal{F}$ denote the set of possible features (e.g., specific keywords, sentiment, verb tense, demographic information, etc.) that the model can be instructed to focus on or ignore when generating responses.
We consider a set of natural language instructions to focus or rule out specified features in $\mathcal{F}$ which we term the focus instruction set $\mathcal{I}_{\text{focus}}$. Explicitly, we define $\mathcal{I}_{\text{focus}}$ as
[…]
$,\;{\color[rgb]{0,0.62109375,0.44921875}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.62109375,0.44921875}\text{focus}(F_{i})\wedge\text{ignore}(F_{j})}\mid F_{i},F_{j}\in\mathcal{F}\},$
where: $\emptyset$ denotes an empty focus instruction with no features to focus on or to ignore; $\text{focus}(F_{i})$ is an instruction to focus on feature $F_{i}$; $\text{ignore}(F_{j})$ is an instruction to ignore feature $F_{j}$; and $\text{focus}(F_{i})\wedge\text{ignore}(F_{j})$ is an instruction to focus on feature $F_{i}$ whilst ignoring feature $F_{j}$. We include the default prompt in order to aid the model in learning the underlying task as well as the ability to refocus its attention on specified features during FIT. For specific examples of focus instructions that we consider, see Appendix E.
Figure 2: Example of focus labels. Focus labels for a modified example from BBQ. Here, age is a spurious feature.
Consider a sample $(x,y)\sim p_{\text{data}}$ drawn from an underlying data distribution, and a focus instruction $I_{\text{focus}}$ drawn from a distribution $p_{\mathcal{I}_{\text{focus}}}$ over the set of focus instructions $\mathcal{I}_{\text{focus}}$. Then the likelihood of response $y$ conditioned on input $x$, task instruction $I$ (as in standard IT), and focus-instruction $I_{\text{focus}}$ is modelled as $p_{\theta}(y|I,I_{\text{focus}},x)$.
FIT Training. Consider a classification task with finite label space $\mathcal{Y}$, where a single core feature $C\in\mathcal{F}$ is fully predictive of label $y\in\mathcal{Y}$ given input $x$ at both training time and under distribution shift (Koh et al., 2021). We also consider spurious features $S\in\mathcal{S}\subseteq\mathcal{F}$ from a subset of spurious features $\mathcal{S}$, where feature values $s\in\text{Val}(S)$ for some spurious feature $S\in\mathcal{S}$ correlate with a label $y_{s}\in\mathcal{Y}$, where this correlation may change under distribution shift (Ming et al., 2022). Finally, we define $\mathcal{F}$ as the set of features that may be included in focus instructions during training, consisting of the core feature and the set of spurious features =$\mathcal{F}=\{C\}\cup\mathcal{S}$.
For a sample $(x,y)\sim p_{\text{data}}$, we specify the focus label $y_{\text{focus}}=y_{\text{focus}}(I_{\text{focus}},s,y)\in\mathcal{Y}$ that depends on the ground truth label $y$, focus instruction $I_{\text{focus}}\in\mathcal{I}_{\text{focus}}$ and spurious feature value $s\in\text{Val}(S)$ present in $x$. Intuitively, we define focus label $y_{\text{focus}}$ as $y_{\text{focus}}=y$ when either no focus features are specified (i.e., using the empty focus instruction), when the focus is on the underlying core feature $C$, or when ignoring a spurious feature $S$; but when either the focus is on a spurious feature or the core feature is ignored, $y_{\text{focus}}$ is defined as $y_{\text{focus}}=y_{s}$, where $y_{s}$ is the label spuriously correlated with the particular spurious feature value $s$ present in input $x$. This changing target $y_{\text{focus}}$ trains the model to learn to adjust its responses based on feature specifications implemented through focus instructions. More formally, we define $y_{\text{focus}}$ as:
$y_{\text{focus}}=\begin{cases}y&\text{if }I_{\text{focus}}\in\mathcal{I}_{\text{focus}}^{C};\\ y_{s}&\text{if }I_{\text{focus}}\in\mathcal{I}_{\text{focus}}^{s}\;,\end{cases}$ (3)
where we define the task-casual and spurious instruction target sets as
$\mathcal{I}_{\text{focus}}^{y}$ =$=\{\emptyset,\;\text{focus}(C),\;\text{focus}(C)\wedge\text{ignore}(S),\;$ (4)
$\qquad\qquad\qquad\qquad\;\;\text{ignore}(S)\mid S\in\mathcal{S}\},$ (5)
$\mathcal{I}_{\text{focus}}^{y_{s}}$ =$=\{\text{focus}(S),\;\text{focus}(S)\wedge\text{ignore}(F_{j}),$ (6)
$\ \qquad\qquad\qquad\qquad\mid F_{j}\in\mathcal{F}\setminus\{S\}\},$ (7)
respectively. See Figure 2 for a concrete example showing the focus label values for an example from the MNLI dataset under different focus instructions.
The objective of FIT training is to minimise the negative log-likelihood (NLL) of the response $y_{\text{focus}}$ conditioned on $I,I_{\text{focus}},x$. Formally, as a form of expected-risk minimisation (ERM) (Vapnik et al., 1998), we define the FT loss objective as:
$\min_{\theta}\mathbb{E}_{(x,y)\sim p_{\text{data}},\;I_{\text{focus}}\sim p_{\mathcal{I}_{\text{focus}}}}\;\left[-\log p_{\theta}\left(y_{\text{focus}}\mid I,I_{\text{focus}},x\right)\right].$ (8)
We define $p_{\mathcal{I}_{\text{focus}}}(\mathcal{I}_{\text{focus}})$ by placing a small probability mass on the empty focus instruction prompt $\emptyset$ in order to aid in learning the underlying task, and then uniformly distribute the remaining probability mass over the remaining non-empty focus instructions. The objective in Equation 8 can be optimised through sampling using stochastic gradient descent (SGD) with popular optimisers such as AdamW (Loshchilov & Hutter, 2019). Further details on FT optimisation are provided in Appendix C.
Evaluating FIT under spurious correlations. After introducing FIT above, we now turn to settings where we can empirically train and evaluate it. A key aspect of our evaluation is the use of known spurious correlations, which simulate real-world scenarios where models can be misled by features that are spuriously predictive of the output label. By adjusting the co-occurrence rate between spurious features and their associated labels, we can test FIT’s ability to dynamically steer a model’s responses depending on the features on which it is focusing or ignoring.
We define the co-ocurrence rate, or predictivity (Hermann et al., 2024), between spurious feature values and the label with which they are spuriously correlated by $\rho_{\text{spurious}}$. Specifically:
Definition 3.1.
(Defining $\rho_{\text{spurious}}$) . Let $S\in\mathcal{S}\subseteq\mathcal{F}$ denote a spurious feature. Suppose that a value of $S$, say $s\in\text{Val}(S)$, is spuriously correlated with label $y_{s}$. Then, for ground truth dataset label $Y$, we define
$\rho_{\text{spurious}}(s)=\mathbb{P}(Y=y_{s}|S=s)$ (9)
By varying $\rho_{\text{spurious}}(s)$, we can control the predictivity of spurious features and observe the model’s behaviour when focusing on or ignoring these features as well as core features under distribution shift.
To examine how well models can utilise focus instructions for steerable behaviour in a controlled environment, we construct synthetic datasets such that the spurious feature $S$ is not predictive of the ground-truth label $Y$ (i.e., $Y\perp\!\!\!\perp S$) and the core feature $C$ is not predictive of the spurious label $Y_{S}$ (i.e., $Y_{S}\perp\!\!\!\perp C$). We enforce these conditions by setting $\rho_{\text{spurious}}=1/N$ (for $N$ label classes) and ensuring a balanced label distribution during training. In Appendix I and Appendix K, we verify that our synthetic SS and SMNLI training sets introduced in Section 4.1 and Figure 4 respectively, satisfy these independence assumptions.
Next, we evaluate FIT across several test sets with varying predictivity levels:
  • $\mathcal{D}_{\text{iid}}$: Held-out test samples with the same $\rho_{\text{spurious}}$ as in the training set.
  • $\mathcal{D}_{\text{high}}$: Test samples with a higher $\rho_{\text{spurious}}$ than in the training set.
  • $\mathcal{D}_{\text{low}}$: Test samples with a lower $\rho_{\text{spurious}}$ than in the training set.
  • $\mathcal{D}_{\text{flipped}}$: Test samples where spurious feature values are flipped to co-occur with different labels than in the training set, with the same high $\rho_{\text{spurious}}$ as in $\mathcal{D}_{\text{high}}$.
We further evaluate FIT under another form distribution shift, specifically on our SMNLI dataset (c.f. Section 4.2).. Here, the specific values taken by spurious features do not overlap between the training and test sets.
  • $\mathcal{D}^{s}_{\cdot}$: Test datasets where the spurious feature values are distinct from those within the training set. Here, we use the same predictivity levels as in the initial datasets presented above.
Note that, while we define FIT with respect to annotated spurious features, this requirement can be alleviated by, e.g., combining FIT with automated spurious feature identification methods (Wang et al., 2022; see Appendix B for further discussion).
4 Experiments
In this section we empirically validate the effectiveness of FIT across a range of popular LLMs of varying sizes and on different NLP datasets, including classification and multi-choice question-answering (MCQA) tasks.
Before reporting the main results, we introduce the evaluation metric (focus accuracy) that we report, baselines, models, and training settings used throughout the experiments. In Section 4.1, we first verify that FIT performs well on the simpler SS dataset, a synthetic sentiment analysis dataset derived from SST-5 (Socher et al., 2013b). We then demonstrate in Section 4.2 that FIT generalises to more complex features and handles feature-value distribution shifts on the SMNLI dataset, a sub-sampled version of the MNLI dataset (Williams et al., 2018). Finally, in Section 4.3, we show that FIT has practical, real-world impact by effectively mitigating bias in the BBQ dataset (Parrish et al., 2022), where we further illustrate FIT’s ability to generalise to new features seen for the first time when performing inference.
Although the primary focus of FIT is on adaptively steering LLMs at inference time, which is what we focus on in this paper, we include an additional debiasing experiment for comparison on the BBQ dataset in Appendix H for completeness. While bias mitigation is a valuable and natural application of FIT, it is not its primary objective. Instead, this inclusion highlights FIT’s broader utility as a tool for model control and adaptability, demonstrating that it performs on par with dedicated bias mitigation techniques while also offering the unique advantage of test-time steerability.
Metrics. We define the focus accuracy for a focus instruction $I_{\text{focus}}\in\mathcal{I}_{\text{focus}}$ as the proportion of samples where the model’s prediction aligns with the focus label, $y_{\text{focus}}$, as specified in Equation 3. Specifically, for each sample $(x,y)\in\mathcal{D}$, the model produces a prediction $\hat{y}\sim p_{\theta}(\cdot\mid I,I_{\text{focus}},x)$ based on a fixed focus instruction $I_{\text{focus}}\in\mathcal{I}_{\text{focus}}$. The focus label, $y_{\text{focus}}=y_{\text{focus}}(I_{\text{focus}},s,y)$, corresponds to the target output given the focus instruction for the input $x$ with ground truth label $y$, a spurious feature value $s$ present in $x$. Focus accuracy for focus instruction $I_{\text{focus}}$, denoted $\mathcal{A}_{\text{focus}}(I_{\text{focus}})$, is computed as the fraction of correct predictions with respect to the focus label:
$\mathcal{A}_{\text{focus}}(I_{\text{focus}})=\frac{1}{|\mathcal{D}|}\sum_{(x,y)\in\mathcal{D}}\mathbf{1}(\hat{y}=y_{\text{focus}}),$ (10)
where $\mathbf{1}(\hat{y}=y_{\text{focus}})$ is the indicator function that equals 1 if the model’s prediction $\hat{y}$ matches the focus label $y_{\text{focus}}$, and 0 otherwise.
We report focus accuracy for each model on all dataset splits, using the prompt types and focus instructions detailed in Appendix E. Generations are evaluated through simple pattern matching due to the use of constrained beam decoding (Anderson et al., 2017). Further details are provided in Appendix D.
Models and training settings. We evaluate FIT using three popular LLMs that span a range of model sizes: Llama-3.1-8B-Instruct (Dubey et al., 2024), Mistral-7B-Instruct-v0.3  (Jiang et al., 2023), and Vicuna-13B-v1.5  (Chiang et al., 2023). The models are fine-tuned using parameter-efficient SFT with LoRA (Hu et al., 2021), leveraging Hugging Face’s SFTTrainer (Wolf et al., 2020) with default hyperparameters. Early stopping is applied based on validation loss, as defined in Equation 8. For generation, we use constrained beam decoding (Anderson et al., 2017) and use fully verbalised (natural language) labels during both training and testing, except for the multi-choice BBQ dataset. For further training details, refer to Appendix C.
Figure 3: SS focus accuracies ($\uparrow$). Focus accuracy ($\mathcal{A}_{\text{focus}}$) of baselines and FIT on the SS dataset.
Baselines. We compare against the following baselines in the main section of the paper: a few-shot baseline (Manikandan et al., 2023) and a SFT baseline. The SFT baseline, $\text{SFT}(y_{\text{focus}})$, follows the same setup as the FIT method (trained on sampled inputs and focus labels), but without the inclusion of focus instructions during training. This ensures a fair comparison between FIT and the baseline, as both methods are trained on the same examples and labels (i.e., focus labels $y_{\text{focus}}$), with the only difference being the inclusion of focus instructions in FIT. This setup allows us to isolate and evaluate the specific impact of incorporating focus instructions in FIT. The few-shot baseline involves using 5 in-context examples uniformly sampled at random from the training set for each test example, where we use the same focus instruction for each in-context sample as for the test sample. In Appendix F, we include two additional baselines: zero-shot and vanilla SFT for a more complete comparison with FIT. Further details of baselines and their results in comparison to FIT can be found in Appendix F.
4.1 Validation of FIT on the SS dataset
Figure 4: SMNLI focus accuracies ($\uparrow$). Focus accuracy ($\mathcal{A}_{\text{focus}}$) of baselines and FIT models on the (a) SMNLI standard test sets $\mathcal{D}$, and (b) SMNLI test sets under feature value shift $\mathcal{D}^{s}$.
[…]
Results. Figure 3 (a) shows the focus accuracy results of the three LLMs on the SS dataset using few-shot prompting, after $\text{SFT}(y_{\text{focus}})$ and after FIT. We see that across all focus instructions and all models, FIT shows significant improvement over the baselines, achieving very high focus accuracy across all focus instruction types and across all test sets with varying predictivity levels.
Key takeaways. High focus accuracy on SS indicates that FIT successfully steers model responses based on the feature on which it is instructed to focus or to not focus on.
4.2 FIT performs well with more complex features on the SMNLI dataset and generalises under distribution shift
[…]
Moreover, for SMNLI, we hold out specific genres at test time to evaluate our model’s ability to generalise under distribution shift when feature values change. We do this by sub-sampling a held-out portion of the MNLI dataset. During training, we use three selected genres (government, fiction, and travel) to evaluate our models. At test time, we additionally add three held-out genres (facetoface, nineeleven, and verbatim). We again ensure that $\rho_{\text{spurious}}$ is constant within each dataset split for all feature values, and use the same set of corresponding $\rho_{\text{spurious}}$ as within the SMNLI test sets described above. Further details of the SMNLI dataset can be found in Appendix K.
Results. Figure 4 (a) depicts the focus accuracy results of the three models on the SMNLI test splits. We observe that for the more complex feature of genre, FIT achieves very high focus accuracy, significantly improving over the baselines. This demonstrates that FIT effectively trains the model to handle more complex features, allowing it to dynamically focus on or disregard these features when making predictions.
Figure 4 (b) shows the focus accuracy of models on the feature-shifted test sets. When focusing on the core feature or ignoring the spurious feature, the model maintains strong performance in terms of focus accuracy, even on unseen genre values (over $80\%$ focus accuracy for FIT models on the third row of Figure 4) (b). Note that, while we observe low focus accuracy when focusing on spurious features, this is expected, as the spurious labels associated with these new genres were not encountered during training. Thus when focusing on these features the model does not know what label to predict. This result highlights that the focus-tuned models have indeed learned spurious associations during training and correctly reproduces them when instructed to focus on these spurious features, even for new spurious feature values. When instructed to focus on the core feature or when instructed to ignore the spurious feature, the model still shows strong generalisation in the presence of distribution shift.
Key takeaways. FIT achieves high focus accuracy on more complex features and maintains strong performance under distribution shift in terms of feature values. This demonstrates FIT’s ability to generalise to new contexts and reliably handle changing feature values.
4.3 FIT steers behaviour in the presence of social bias data and generalises to unseen features
[…]
Figure 5: BBQ focus accuracies ($\uparrow$). Focus accuracy ($\mathcal{A}_{\text{focus}}$) of baselines and FIT on the BBQ dataset.
Results. Figure 5 shows the focus accuracy results of the three models on the BBQ dataset, visualising performance on features seen during training and unseen, held-out features. The models demonstrate high and comparable focus accuracy across both seen and unseen bias features, indicating that FIT generalises well to unseen features, including nuanced reasoning about group stereotypes. This highlights the usefulness of FIT in mitigating social biases in LLM responses. Specifically, FIT can effectively learn, reason about, and rule out biases when formulating responses, making it a practical tool for bias mitigation.
Key takeaways. FIT can effectively teach models to adjust their responses based on knowledge of social biases. This ability generalises to biases not seen during FIT training, indicating FIT’s utility for bias mitigation.
5 Ablation
Generalisation to different test-time prompt formats. As observed in the IT literature, instruction-tuned models sometimes memorise instruction formats and struggle to follow paraphrased instructions at test time (Ghosh et al., 2024). In Appendix G (Figure 9), we compare the performance of models on the SMNLI dataset when using the same focus instructions at training and test time versus using paraphrased instructions at test time. We generate 10 different test-time focus instructions of each instruction type defined in Equation 2 by paraphrasing the existing focus instruction using ChatGPT (OpenAI, 2022). The results show minimal variation in focus accuracy across different dataset splits and focus features, even when testing on paraphrased prompts, indicating that FIT indeed teaches models a general capacity to focus on or ignore features regardless of the specific way that focus instructions are phrased.
Instruction Following After FIT. Prior studies suggest that SFT can impair the instruction-following abilities of LLMs (Fu et al., 2024; Dou et al., 2024). We evaluate whether FIT impacts instruction adherence by comparing pre-trained and FIT models trained on the SMNLI dataset (see Section 4.2). Using 500 samples from the Alpaca-GPT instruction-tuning dataset (Peng et al., 2023), responses were rated by GPT-4o (Achiam et al., 2023) on a 1–5 scale, where 5 denotes perfect alignment and 1 indicates a completely incorrect response.
LLaMA Mistral Vicuna
Pre-Trained Avg. Rating 3.51 3.65 3.46
FIT Avg. Rating 3.45 3.65 3.50
$p$-value $\bm{0.57}_{>0.05}$ $\bm{0.81}_{>0.05}$ $\bm{0.41}_{>0.05}$
Table 1: Instruction-Following Results. For each model (columns), we report the pre-trained and FIT average GPT-40 ratings, and the two-sided Wilcoxon Signed-Rank $p$-value testing the difference between the distributions of ratings.
To test for significant differences in instruction-following performance, we conducted a two-sided Wilcoxon Signed-Rank Test (Wilcoxon, 1992) on paired ratings. The null hypothesis assumes no difference in the median ratings between pre-trained and FIT models. The results, summarized in Table 1, show no statistically significant differences ($p>0.05$) in instruction-following performance across all models. These findings demonstrate that FIT maintains general instruction-following capabilities whilst additionally enhancing test-time steerability.
6 Conclusion
In this work, we introduce Focus Instruction Tuning (FIT), a method designed to steer the behaviour of LLMs by focusing on or ignoring specific features when formulating responses. Across a range of tasks and settings, we demonstrate that FIT provides dynamic and precise control over LLM behaviour at inference time, enabling users to adapt model responses even in the context of distribution shifts over feature values or when generalising to unseen features. Furthermore, our approach can address challenges such as mitigating the influence of known stereotypes that might otherwise impact responses, showcasing one of its many applications. Thus, FIT represents a step toward enabling more robust, steerable, fair, and controllable LLMs.
Appendix A Impact Statement
The ability to dynamically steer model behaviour by focusing on or ignoring features, as enabled by FIT, holds significant potential for reducing algorithmic discrimination and mitigating harms. Practitioners can leverage FIT to identify and correct biases by measuring discrepancies in behaviour when a model focuses on or ignores specific features. Additionally, FIT enhances explainability by attributing model predictions to input features, enabling more transparent and productive human-AI collaboration. This supports ethical and responsible decision-making by assessing whether predictions are justified. FIT also enhances robustness by prioritising stable core features expected to generalise across domains while ignoring spurious, domain-specific biases, making it a valuable tool for fairness, explainability, and robustness. However, risks include potential misuse by bad actors to bias models, though this is not unique to FIT and could already be achieved through biased fine-tuning.
Appendix B Limitations and Future Work
Requirement for annotated spurious features. While FIT relies on prior identification of spurious features and their focus labels, this requirement does not limit its practical applicability. Instead, it reflects standard industry and research practices for constructing transparent and reliable models. Below, we clarify how FIT remains adaptive and versatile even when feature annotation is partial or evolving:
  • Alignment with Established Practices: FIT’s reliance on pre-identified spurious features aligns with widely adopted industry and research norms (OpenAI, 2024; Microsoft, 2020). Identifying potential spurious features and confounders in datasets is a foundational step in achieving robust machine learning systems. This process ensures that both training and validation phases are informed by an understanding of data correlations, minimising the risk of deploying models with unknown biases.
  • Regulatory and Ethical Expectations: Regulatory frameworks and ethical guidelines increasingly require the explicit identification and mitigation of problematic features (of the European Parliament, 2016) Corresponding initiatives aim to define and enforce measurable categories of “violating behaviour” in AI models. By providing a mechanism to steer model behaviour based on these identified features, FIT effectively complements efforts to promote fair and transparent predictions (Guldimann et al., 2024; Zeng et al., 2024).
  • Post-Deployment Mitigation: Despite careful pre-deployment analysis, spurious features or correlations may only become apparent once a model is in active use. FIT accommodates this by allowing developers to incorporate newly identified spurious features via updated focus instructions, enabling rapid iterative refinement without retraining from scratch. This adaptability ensures continuous improvement, even in highly dynamic environments.
  • FIT’s Versatility Without Exhaustive Pre-Identification: Crucially, FIT does not require an exhaustive list of spurious features to be effective. For instance, a user can provide focus instructions such as “focus on casual” without enumerating every possible irrelevant attribute in the dataset. This flexibility expands FIT’s applicability to scenarios where feature annotation is incomplete or ongoing.
  • Compatibility with Automated Spurious Feature Identification: FIT also works seamlessly with automated methods for detecting spurious features (Wang & Culotta, 2020; Wang et al., 2022; Zhou et al., 2023; Zheng et al., 2024). Whether spurious features are labelled manually or derived from algorithmic detection, they can be harnessed by FIT’s focus instructions at inference time. This compatibility enables a comprehensive approach to managing known issues and responding to newly uncovered features as they arise.
In summary, annotating spurious features beforehand is not a strict limitation. FIT can be flexibly applied, allowing model behaviour to evolve in tandem with new feature discoveries or changing requirements, making it a broadly applicable technique for steering model outputs based on both prior knowledge and ongoing insights.
Scope of Experiments and Extensions to Open-Ended Tasks. Our experiments primarily focus on classification and multiple-choice QA datasets due to the cost and challenges associated with curating high-quality datasets for open-ended NLG tasks. However, this reflects a pragmatic prioritisation of introducing a novel methodology over exhaustive data collection, rather than a limitation of FIT itself. Extending FIT to open-ended tasks, such as summarisation or translation, remains an exciting direction for future research, as does exploring its ability to generalise across diverse task categories using setups similar to FLAN (Longpre et al., 2023).
Overlapping Features and Ambiguities. Additionally, our evaluation on the HANS dataset Appendix L revealed challenges when addressing overlapping or less-distinctive features. While FIT demonstrated strong performance in generalising and steering models based on identified features, overlapping heuristics can introduce ambiguity, highlighting the need for further refinements in handling such cases. Despite these limitations, FIT represents a promising foundation for enabling more robust, fair, and controllable LLMs across a range of tasks.
Appendix C FT Training and Optimisation Settings
FT Optimisation. Algorithm 1 gives precise details on how we implement FIT in practice when performing ERM of a model on a given training set. In particular, it shows how we approach optimising the FIT training objective given in Equation 8.
Algorithm 1 Algorithm for Focus Instruction Tuning (FIT) Training Procedure to Optimise Equation 8.
1:  Input: Dataset $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$, The feature set contains $\mathcal{F}$, instruction $I$, model parameters $\theta$, batch size $B$, number of epochs $E$, step size $\eta$, and mapping function $y_{\text{focus}}=y_{\text{focus}}(I_{\text{focus}},y,s)$.
2:  Initialise: Model parameters $\theta$, optimiser
3:  for epoch $=1$ to $E$ do
4:     for mini-batch $\{(x^{b},y^{b})\}_{b=1}^{B}$ from $\mathcal{D}$ do
5:        for each $(x^{b},y^{b})$ in the mini-batch do
6:           Identify spurious feature value $s^{b}$ in $x^{b}$.
7:           Sample focus instruction $I_{\text{focus}}^{b}\sim p_{\mathcal{I}_{\text{focus}}}$,
8:           Compute $y_{\text{focus}}^{b}=y_{\text{focus}}(I_{\text{focus}}^{b},s^{b},y^{b})$
9:        end for
10:        Compute average loss given through empirical estimator of the loss defined in Equation 8 over the batch:
$\ell(\theta)=\frac{1}{B}\sum_{b=1}^{B}-\log p_{\theta}(y_{\text{focus}}^{b}|I,I_{\text{focus}}^{b},x^{b})$
11:        Update model parameters $\theta$ using optimiser:
$\theta\leftarrow\theta-\eta\nabla_{\theta}\ell(\theta)$
12:     end for
13:  end for
14:  Output: Optimised model parameters $\theta$
FT training settings. We use the SFTTrainer class from HuggingFace (Wolf et al., 2020) and use all of the default training settings for performing SFT of LLMs. Furthermore, we define $p(\mathcal{I}_{\text{focus}})$ by placing a small probability (in our experiments, 0.05) on the empty focus instruction $\emptyset$. We then uniformly distribute the remaining probability mass over the non-empty focus instructions.
We implement early stopping on a held-out validation set based on the cross-entropy loss over focus labels $y_{\text{focus}}$ corresponding to randomly sampled focus instructions - this matches the context in which the models will be evaluated. We obtain this set by splitting our training set in a 90/10% ratio for training and validation splits respectively. We use a patience of $4$ validation evaluation steps, which occur after a fixed number of steps.
We use LoRA (Hu et al., 2021) for parameter-efficient fine-tuning. We target the query and value projection matrices within each LLM and use LoRA $r=16$ and $\alpha=32$ across models.
Choice of $\rho_{\text{spurious}}$ during training. In our synthetic experiments, we set up a controlled environment by imposing two independence conditions: $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$. These ensure that (i) the ground-truth label cannot be predicted using the spurious feature $S$, and (ii) the spurious label cannot be predicted using the core feature $C$. By removing direct correlations between these features and labels, the model is leans to focus on the specified feature, without being influenced by the other feature, avoiding any potential shortcuts that could be exploited if these conditions did not hold.
  • Independence $Y\perp\!\!\!\perp S$: This condition prevents the model from leveraging spurious feature $S$ to predict ground-truth label $Y$. With no predictive signal from $S$ to $Y$, the model must rely exclusively on the core feature $C$ for accurate label predictions. This design choice safeguards the model from overfitting to spurious correlations, thereby maintaining robust performance under distribution shifts. Moreover, removing any inherent relationship between $S$ and $Y$ ensures that for focus instruction intending for the model to utilise the core feature $C$ only during inference, the model cannot exploit a potential shortcut using $S$; it must utilise the core feature alone for prediction in this scenario enabling prediction only through the specified feature indicated through the focus instruction passed to the model.
  • Independence $Y_{S}\perp\!\!\!\perp C$: This condition serves a complementary role in preventing the model from exploiting the core feature $C$ when predicting spurious labels $Y_{S}$. By ensuring $C$ carries no information about $Y_{S}$, the model cannot use the true task feature $C$ as a shortcut for spurious-label predictions; it must again learn to only use the specified feature within the passed focus instructions alone for making predictions.
While these conditions represent an ideal setting, they are not strictly necessary for FIT to work in practice. Indeed, real-world data rarely satisfies such perfect independence, and we illustrate the robustness of the method in more realistic scenarios through our BBQ experiments in Section 4.3 where correlations between $Y$ and $S$ or between $C$ and $Y_{S}$ may exist as no subsampling or dataset manipulations have been made. By examining both the controlled environment and more naturalistic datasets, we demonstrate that our approach can handle scenarios with varying degrees of spurious correlations.
To achieve this independence in our synthetic SS and SMNLI datasets, we set $\rho_{\text{spurious}}=1/N$, where $N$ is the number of class labels. Additionally, we enforce a balanced label distribution in the training set to eliminate any indirect biases that could correlate $S$ with $Y$. As shown in Appendix I and Appendix K, these conditions are sufficient to guarantee $Y\perp\!\!\!\perp S$ in the training data, enabling the model to effectively learn steerable behaviour from focus instructions.
Appendix D Evaluation Metrics
Generation settings. We generate responses from our FT model using constrained beam-decoding (Anderson et al., 2017) with $8$ beams. This ensures that the answer labels for each classification task that we investigate appear in the model’s output. We limit the maximum number of newly generated tokens to be $5$ to stop any unnecessary text given after the model’s initial classification prediction.
Computing the focus accuracy metric. We report the focus accuracy $\mathcal{A}_{\text{focus}}$ of generations when evaluating FT models. As we are guaranteed to include the task labels within the model’s response through constrained decoding, we simply check to see if the focus label, $y_{\text{focus}}$, is within the model’s response or not in order to determine if the model’s response is correct.
Appendix E FIT Focus Instructions and Prompt Templates
Prompt templates. Here we provide the prompt templates that we use for FIT. Note, that we drop the feature considerations section and focus instruction from the prompt to form the instruction prompt $\emptyset$ with no focus instruction.
SS Focus instruction prompt templates $\mathcal{I}_{\text{focus}}$ <INSTRUCTION>
You are a language model performing sentiment analysis on a binary dataset, making predictions from the labels [negative, positive]. Make your prediction based on the relevant features described below, focusing on the specified features and ignoring those deemed irrelevant. For the input below, output either negative or positive ONLY for your prediction of the input’s label.
<END OF INSTRUCTION>
<FEATURE CONSIDERATIONS>
{focus instruction - $I_{\text{focus}}$}
<END FEATURE CONSIDERATIONS>
### Input: {input}
### Output:
NLI focus instruction prompt template $\mathcal{I}_{\text{focus}}$ <INSTRUCTION>
You are a language model performing natural language inference (NLI) on a dataset with three labels [entailment, neutral, contradiction]. For the inputs consisting of premise and hypothesis below, make your prediction based on the relevant features described, focusing on the specified features and ignoring those deemed irrelevant. Output either entailment, neutral, or contradiction ONLY for your prediction of the input’s label.
<END OF INSTRUCTION>
<FEATURE CONSIDERATIONS>
{focus instruction - $I_{\text{focus}}$}
<END FEATURE CONSIDERATIONS>
### Premise: {premise}
### Hypothesis: {hypothesis}
### Output:
BBQ focus instruction prompt template $\mathcal{I}_{\text{focus}}$ <INSTRUCTION>
You are a language model answering a QUESTION based on a provided CONTEXT, choosing to ANSWER with one of three options: 1, 2, or 3. Each option corresponds to an answer from the provided ANSWER OPTIONS. You must output EXACTLY one of these numbers (1, 2, or 3) representing the answer you select.
<END OF INSTRUCTION>
<FEATURE CONSIDERATIONS>
{focus instruction - $I_{\text{focus}}$}
<END FEATURE CONSIDERATIONS>
### CONTEXT: {context}
### QUESTION: {question}
### ANSWER OPTIONS:
(1) {ans0}
(2) {ans1} (3) {ans2} ### ANSWER:
Focus instructions. We consider the following focus instruction formats for the different focus instructions introduced in Equation 2 which are used for FIT training and evaluation:
Focus instructions $\mathcal{I}_{\text{focus}}$ For features $F_{i},F_{j}\in\mathcal{F}$:
Focus instructions $\text{focus}(F_{i})$:
Direct your attention solely to $F_{i}$. Concentrate all your reasoning on $F_{i}$. Make $F_{i}$ the central factor in your decision. Base your judgment exclusively on $F_{i}$. Pay attention only to $F_{i}$ when making your prediction. Use $F_{i}$ as the key input for your evaluation. Focus entirely on $F_{i}$ and ignore other aspects. Rely exclusively on $F_{i}$ to reach your conclusion. Consider only $F_{i}$ and disregard all else. Let $F_{i}$ be the primary basis for your decision.
Ignore instructions $\text{ignore}(F_{i})$: Completely rule out $F_{i}$ from your reasoning. Disregard any influence of $F_{i}$ in your prediction. Treat $F_{i}$ as irrelevant to your decision-making process. Exclude $F_{i}$ entirely from your evaluation. Do not let $F_{i}$ play any role in your assessment. Intervene to prevent $F_{i}$ from affecting your prediction. Ensure that $F_{i}$ has no bearing on your final decision. Block $F_{i}$ from contributing to your reasoning. Negate the impact of $F_{i}$ in your prediction. Ruling out $F_{i}$ is crucial—do not let it affect your decision.
Focus and Ignore instructions $\text{focus}(F_{i})\wedge\text{ignore}(F_{j})$ Focus specifically on $F_{i}$. Disregard $F_{j}$ in your decision-making process. Base your prediction solely on $F_{i}$. Exclude $F_{j}$. Direct all your attention to $F_{i}$. Block out $F_{j}$ from your prediction. Consider only $F_{i}$ in your reasoning. Rule out $F_{j}$ in your decision-making. Prioritize $F_{i}$. Completely ignore $F_{j}$ in your prediction. Do not consider $F_{j}$ in your decision-making process. Focus exclusively on $F_{i}$. Ignore any influence of $F_{j}$. Concentrate on $F_{i}$ in your prediction. Disregard $F_{j}$ entirely. Base your analysis solely on $F_{i}$. Rule out $F_{j}$ in your prediction. Shift your focus to $F_{i}$. Do not pay attention to $F_{j}$ in your decision-making process. Rely only on $F_{i}$.
Appendix F Additional Baselines Results
In addition to the baselines that we present in the main paper, Few-shot and $\text{SFT}(y_{\text{focus}})$, we include two additional baselines to further supplement these results. We give the complete list of baselines that we consider below:
Zero-shot baseline. Finally, we include a zero-shot inference baseline using the original pre-trained models without additional fine-tuning on our spurious datasets. No in-context examples are used at inference time, and the model is not trained at all beyond it’s pre-training. The model is tested on the full set of focus instructions prompts detailed in Equation 2.
Few-shot baseline. This second baseline compares FIT training to few-shot inference using the original pre-trained models without additional fine-tuning on our spurious datasets. Specifically, we use 5 in-context examples across all datasets. For the in-context examples, we concatenate multiple examples one after the other, including the instructional prompt only for the first in-context example and the final test example. Each in-context example contains the same focus instruction as the test example for which they serve as context. The model is tested on the full set of focus instructions prompts detailed in Equation 2.
SFT($y_{\text{focus}}$) baseline. We implement an SFT baseline that follows the same training procedure as FIT, except during training, we exclude any focus instructions from the input prompts while still training on the focus labels. This provides a fair comparison with FIT, as the models are trained on the same input text and label pairs. The rest of the training setup, including hyperparameters and early stopping, remains identical to the FIT training setup. The model is tested on the full set of focus instructions prompts detailed in Equation 2.
SFT($y$) baseline. We implement a vanilla SFT baseline that simply trains a model using SFT on inputs and their ground truth labels (as opposed to focus labels in the SFT($y_{\text{focus}}$) baseline). During training, only standard IT prompts are used, with no additional focus instructions included. The rest of the training setup, including hyperparameters and early stopping, remains identical to the FIT training setup. The model is tested on the full set of focus instructions prompts detailed in Equation 2.
We give the full set of results for all datasets and models across the complete set of baselines listed above in Figure 6, Figure 7, and Figure 8.
Figure 6: Full Baseline vs FIT Focus accuracy ($\uparrow$) on SS. Figure giving focus accuracies ($\mathcal{A}_{\text{focus}}$) of the additional baselines compared to the focus accuracy of FIT on the SS dataset.
Figure 7: Full Baseline vs FIT Focus accuracy ($\uparrow$) on SMNLI. Figure giving focus accuracies ($\mathcal{A}_{\text{focus}}$) of the additional baselines compared to the focus accuracy of FIT on the SMNLI dataset.
Figure 8: Full Baseline vs FIT Focus accuracy ($\uparrow$) on BBQ. Figure giving focus accuracies ($\mathcal{A}_{\text{focus}}$) of the additional baselines compared to the focus accuracy of FIT on the BBQ dataset.
Appendix G SMNLI Ablation of Training and Test time Focus Instruction Rephrasing Differences
We analyse the impact of using the same versus different sets of focus instructions at training and test time when applying FIT models. Specifically, we generate alternative test set focus instructions by paraphrasing the training focus instructions, as shown in Appendix E, using ChatGPT.
Figure 9: Focus accuracy ($\uparrow$) for different training and test focus instruction sets. Figure comparing focus accuracies ($\mathcal{A}_{\text{focus}}$) of sampling from the same (top) and different (bottom) sets of focus instructions at training and test time of models on the SMNLI dataset.
As depicted in Figure 9, the results of this ablation reveal negligible differences between using the same or different focus instruction phrasings during training and testing. This indicates that FIT effectively trains the model to focus on or ignore features, regardless of how the instructions are phrased.
Appendix H Comparison of FIT against a specific debiasing technique
FIT is a general framework designed to enable users to steer a model’s behavior based on specified features. This approach provides enhanced control over model outputs during inference, adding a critical layer of explainability and controllability to model predictions.
While understanding and mitigating biases or spurious correlations is a valuable and natural application of FIT, it is not the sole objective. The broader goal of steerability includes addressing challenges in managing and aligning model behaviour across diverse contexts. For instance, maintaining controllability is crucial in addressing safety alignment fragility, which can emerge after fine-tuning (Bhattacharjee et al., 2024). In such cases, the ability to adapt model responses to align with user specifications ensures safe and reliable deployment.
Experiment. To explore FIT’s broader applicability, we compare its performance as a debiasing method against a well-known debiasing technique: the Product of Experts (PoE) method (Mahabadi et al., 2020). PoE involves training a bias model $f_{B}$, which is trained exclusively on bias features. This bias model mediates the training of the final model $f$ by combining their predictions through an elementwise product: $\sigma(f(x))\odot\sigma(f_{B}(x_{B}))$, where $x\in\mathcal{D}$, for dataset $\mathcal{D}$, and $x_{B}$ represents the biased feature of $x$.
We adapted this approach to our setting by training a bias model on the stereotypical labels within the BBQ dataset. These labels correspond to group-stereotypical associations. For autoregressive models, we further modified the PoE method by extracting and normalising the logits of the first newly generated token position over the set of single tokens representing the answer options.
Results. The results of the debiasing experiment comparing FIT to the PoE method is shown in Figure 10. FIT performs equally as well as the PoE method as shown by the comparing the default prompt accuracy ($\emptyset$) for the PoE models against the $\text{focus}(C)$ results for the FIT models; both metrics correspond to causal accuracy for these prompt types. Indicating that FIT performs just as well as a dedicated debiasing technique.
However, the PoE method requires training two separate models and does not provide steerability at test time as shown by the low focus accuracy on $\text{focus}(S)$. Indeed the model defaults to the ground truth label across all prompt types and does not change behaviour despite different different focus specifications. This highlights the flexibility of FIT, which not only debiases effectively but also enables additional controllability during inference.
Figure 10: Focus Accuracy ($\uparrow$) of FIT against PoE Debiasing Technique. Figure showing the focus accuracies ($\mathcal{A}_{\text{focus}}$) of FIT (bottom row) and the dedicated debaiasing technique, PoE (top row), on the BBQ dataset.
Appendix I Spurious Sentiment (SS) Dataset
We take a pre-existing dataset, in this case SST-5 (Socher et al., 2013a), and modify it in order to induce a known spurious feature and create a spurious binary sentiment analysis dataset.
Data-generating process (DGP). We frame our DGP using a graphical model to describe the synthetic dataset that we create. We follow a similar model to that described in (Arjovsky et al., 2019), specifically the model used for generating their coloured MNIST dataset. We use the following variables within our graphical model:
  • $C$ true underlying sentiment, the core feature within this task, sampled from the original dataset.
  • $\tilde{S}$ proposed spurious feature sample, here this is the presence of the keywords Bayesian or Pineapple. We represent this as a binary vector $S\in\{0,1\}^{2}$, where the first and second components of this vector denote the presence of either the keyword Pineapple or Bayesian respectively. We restrict to consider only one keyword appearing in a given text at a time so that $\text{Val}(S)=\{(1,0),(0,1)\}$.
  • $S$ final spurious feature that is naturally inserted using a LLM into the final SS dataset example $X$. $S$ is a randomly flipped version of the proposed spurious feature $S$.
  • $\tilde{X}$ is a sampled example from the original dataset that we are modifying to inject known spurious correlations.
  • $X$ original example $\tilde{X}$ but augmented to include the spurious feature.
  • $Y$ final label for element $X$.
Figure 11: SS DGP. Graphical model showing the data generating process for modifying examples from the SST-5 dataset to introduce a new spurious keyword feature $S$.
The graphical model describing the DGP of the SS dataset is given in Figure 11. This admits a functional representation in the form:
$C$ $=f_{C}(U_{C});$ (11)
$\tilde{X}$ $=f(C,U_{\tilde{X}});$ (12)
$\tilde{S}$ $=f_{\tilde{S}}(C,U_{\tilde{S}});$ (13)
$S$ $=f_{S}(\tilde{S},U_{S});$ (14)
$X$ $=f_{X}(\tilde{X},S,U_{X});$ (15)
$Y$ $=f_{Y}(C,U_{Y}).$ (16)
where $U_{(\cdot)}$ are variables introducing sources of randomness into the generating process. More explicitly, we consider the following set of equations, where $\mathcal{D}$ denotes the underlying dataset that we are manipulating:
$C\sim\texttt{Ber}(\rho_{C}),\text{ where }\rho_{C}=\rho_{C}(\mathcal{D});$ (17)
$\tilde{X}\sim p_{\mathcal{D}}(\cdot|C)\;,$ (18)
$\tilde{S}=\left(\bm{1}_{C=0},\bm{1}_{C=1}\right);$ (19)
$U_{S}\sim\texttt{Ber}(\rho_{\text{spurious}});$ (20)
$S=U_{S}\tilde{S}+(1-U_{S})(1-\tilde{S});$ (21)
$U_{\text{incld.}}\sim\texttt{Ber}(\rho_{\text{incld.}});$ (22)
$X=\texttt{LLM}(\tilde{X},S);$ (23)
$Y=C,$ (24)
The variable $\rho_{C}$ gives the distribution of sentiment labels in the original binarised SST-5 dataset. Moreover, $p_{\mathcal{D}}(\tilde{x}|C)$ denotes the conditional dataset distribution of the different input texts give $C$ (here we assume that we are just uniformly sampling text with the given sentiment $C$) and $\bm{1}_{(\cdot)}$ denotes the indicator function.
Finally, we prove that $\rho_{\text{spurious}}$ gives the coocurrence rate/predictivity between the label $Y$ and the spurious feature $S$, and is well-defined notation in the sense that it corresponds to Definition 3.1 so that $\rho_{\text{spurious}}(s)=\mathbb{P}(Y=y_{s}|S=s)$, where $y_{s}$ is the label that spurious feature value $s$ is spuriously correlated with.
Proposition I.1.
From the SCEs described above, assuming that we have a balanced label distribution, that is $\rho_{C}=1/2$, we have that
$\mathbb{P}(Y=y_{s}\mid S=s)=\rho_{\text{spurious}}\;.$ (25)
for all spurious feature values $s$.
Proof.
First note that we have by Equation 17 and the assumption that $\rho_{C}=1/2$, that $\mathbb{P}(Y=1)=\mathbb{P}(Y=0)=1/2$. Moreover, using the partition theorem, we have that
$\mathbb{P}(S=s)$ $=\mathbb{P}(S=s\mid\tilde{S}=s)\mathbb{P}(\tilde{S}=s)+\mathbb{P}(S=s\mid\tilde{S}\neq s)\mathbb{P}(\tilde{S}\neq s)$ (26)
$=\rho_{\text{spurious}}\cdot\frac{1}{2}+(1-\rho_{\text{spurious}})\cdot\frac{1}{2}$ (27)
$=\frac{1}{2}.$ (28)
Finally note that $\mathbb{P}(S=s\mid Y=y_{s})=\rho_{\text{spurious}}$ as $Y=y_{s}$ forces $C=y_{s}$, and therefore that $\tilde{S}=s$ by construction.
Putting this all together and utilising Bayes’ rule gives
$\mathbb{P}(Y=y_{s}\mid S=s)$ $=\frac{\mathbb{P}(S=s\mid Y=y_{s})\mathbb{P}(Y=y_{s})}{\mathbb{P}(S=s)}$ (29)
$=\frac{\mathbb{P}(S=s\mid Y=y_{s})\mathbb{P}(Y=y_{s})}{\mathbb{P}(S=s\mid\tilde{S}=s)\mathbb{P}(\tilde{S}=s)+\mathbb{P}(S=s\mid\tilde{S}\neq s)\mathbb{P}(\tilde{S}\neq s)}$ (30)
$=\frac{\rho_{\text{spurious}}\cdot\frac{1}{2}}{\frac{1}{2}}$ (31)
$=\rho_{\text{spurious}}.$ (32)
which gives the result. ∎
Within our experiments in Section 4.1, we always force the label distribution to be balanced, that is $\rho_{C}=1/2$, and assume that within each dataset split, $\rho_{\text{spurious}}$ is the same rate for all spurious feature values. Independence conditions during training for FIT. As specified in Appendix C, we would like to have that $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ during training so that models trained via FIT can effectively learn to leverage focus instructions to make predictions based on specified features. Here, $Y_{S}$ is the spurious label spuriously correlated to spurious feature value $S$. The results below give sufficient conditions for these independence conditions to be specified with respect to the DGP described in Figure 11.
Proposition I.2.
Assuming the SCEs given above and the corresponding DGP described in Figure 11, if we have a uniform label distribution $p(Y=y)$, that is $\rho_{C}=1/2$, and have that $\rho_{\text{spurious}}=1/2$ in the SS training set, then we have that $Y\perp\!\!\!\perp S$.
Proof.
[…]
Proposition I.3.
Assuming the SCEs given above and the corresponding DGP described in Figure 11, if we have a uniform label distribution $p(Y=y)$, that is $\rho_{C}=1/2$, and have that $\rho_{\text{spurious}}=1/2$ in the SS training set, then we have that $Y_{S}\perp\!\!\!\perp C$.
Proof.
[…]
$\mathbb{P}(S=s\mid C=c)$ $=\sum_{\tilde{s}}\mathbb{P}(S=s\mid C=c,\tilde{S}=\tilde{s})\mathbb{P}(\tilde{S}=\tilde{s}\mid C=c)$ (33)
$=\sum_{\tilde{s}}\mathbb{P}(S=s\mid\tilde{S}=\tilde{s})\mathbb{P}(\tilde{S}=\tilde{s}\mid C=c)$ (34)
$=\underbrace{\mathbb{P}(S=s\mid\tilde{S}=s)}_{=\rho_{\text{spurious}}}\underbrace{\mathbb{P}(\tilde{S}=s\mid C=c)}_{=\frac{1}{2}}$ (35)
$+\underbrace{\mathbb{P}(S=s\mid\tilde{S}\neq s)}_{=1-\rho_{\text{spurious}}}\underbrace{\mathbb{P}(\tilde{S}\neq s\mid C=c)}_{\frac{1}{2}}$ (36)
$=\frac{1}{2}.$ (37)
[…]
SCM from this DGP. Through the above data generation process, we introduce a new spurious feature within the dataset $S$, the presence of the keywords Bayesian and Pineapple. Recalling that $S=(1,0)$ and $S=(0,1)$ correspond to insertion of the keywords Pineapple and Bayesian respectively, we introduce the following spurious correlations between feature values of $S$ and label $Y$:
  1. 1.
    The presence of the word Pineapple in the text $X$ is spuriously correlated the label 00 (negative sentiment).
  2. 2.
    The presence of the word Bayesian in the text $X$ is spuriously correlated with the label $1$ (positive sentiment).
The sentiment feature still remains core within the augmented SS dataset, fully predicting the label $Y$ for each dataset example.
The above DGP, through the introduction of spurious feature $S$, induces a SCM that describes the spurious correlation between spurious feature $S$ and the label $Y$. The SCM, shown in Figure 12, follows the style-content decomposition described in (Kaddour et al., 2022), where $U$ is some hidden confounding variable.
Data generation methodology. We use Llama-3.1-70B-Instruct to generate modifications $X$ of original dataset examples $\tilde{X}$ to create new text which include the new keywords feature. The prompt we use for generation when modifying examples to include spurious features is give as:
Figure 12: SS SCM. SCM showing showing the spurious correlation present between the keyword feature $S$ and the label $Y$ of examples within the SS dataset, induced through the described data augmentation process.
Data augmentation prompt You are a language model designed to modify a piece of text to include an additional feature in a simple, natural way while keeping your output as similar as possible to the original text.
Features pineapple: Include the word ‘pineapple’. Bayesian: Include the word ‘Bayesian’. Instructions Ensure the output is grammatically correct. Keep the output as similar as possible to the original text. Make the minimal number of modifications and add the fewest new tokens possible to satisfy the chosen feature. Do not change the sentiment of the original text. Do not significantly alter the length of the output. Incorporate the feature naturally within the original text so that it blends seamlessly with the text’s context. Do not only append additional clauses at the end of the text to include the feature. Inclusions should be case sensitive, e.g., include ‘Bayesian’ BUT NOT ‘bayesian’. Output Only return the modified text, with no additional explanations or reasoning. Should strictly follow the feature description and the set of instructions. Only include the one feature given; the other features SHOULD NOT be included even accidentally.
Appendix J Spurious NLI dataset (SMNLI)
We generate a tertiary NLI dataset, SMNLI, with a known spurious feature. We do this considering the MNLI dataset (Williams et al., 2018). This is a NLI dataset with three labels: entailment (0), neutral (1) and contradiction (2), where data is sampled from 5 underlying categories or genres (telephone, government, travel, fiction or slate). We aim to induce spurious correlations between the underlying genres and labels.
Data-generating process (DGP). We consider a graphical model to describe the DGP of examples within the SMNLI dataset. We use the following variables within our DGP:
  • $C$ NLI relationship between a premise and hypothesis pair, the core feature within this task, sampled from the original dataset.
  • $S$ spurious feature, here this is the genre of the premise and hypothesis. This is a categorical variable.
  • $X$ example from the MNLI dataset.
  • $Y$ final label for element $X$.
Figure 13: SMNLI DGP. Graphical model showing the data generating process for modifying examples from the MNLI dataset to introduce a new spurious keyword feature $S$.
The graphical model described by the DGP for producing the SMNLI dataset is given in Figure 13. Once again, this graphical model can be represented functionally as:
$S$ $=f_{S}(U_{S});$ (38)
$C$ $=f_{C}(S,U_{C});$ (39)
$X$ $=f_{X}(C,E,U_{X});$ (40)
$Y$ $=f_{Y}(C,U_{Y}).$ (41)
More specifically, given the orignal dataset $\mathcal{D}$ that we are sub-sampling from, the functions that we use within the DGP for the SMNLI dataset are given by:
$S$ $\sim\texttt{Cat}(\mathcal{S}),$ (42)
$U_{C}$ $\sim\texttt{Ber}(\rho_{\text{spurious}});$ (43)
$C$ =$=\begin{cases}y_{S}&\text{ if }U_{C}=1;\\ \sim\texttt{Unf}(\text{Val}(Y)\setminus\{y_{S}\})&\text{ if }U_{C}=0;\end{cases};$ (44)
$X$ $\ \sim p_{\mathcal{D}}(\cdot|C,S)$ (45)
$Y$ $=C.$ (46)
Here, $\texttt{Cat}(\mathcal{S})$ is a uniform categorical distribution over spurious feature values (here the underlying genre of the premise-hypothesis pairs). Furthermore, we define $y_{S}$ to be the NLI label that a particular value of $S$ is spuriously correlated with by design. Moreover, $p_{\mathcal{D}}(x|C,S)$ is the conditional distribution over the dataset examples (premise-hypothesis pairs) that have NLI relationship $C$ and genre $S$.
We restrict the genres within the model to $S\in\{\textit{slate},\textit{government},\textit{fiction},\textit{travel}\}$, a subset of the genres of the training set. When creating a distribution shifted test set, we restrict the genres to $S\in\{\textit{facetoface},\textit{nineeleven},\textit{verbatim}\}$. The specific spurious correlations between a genre $s$ and a label $y_{s}$ are chosen to be: $y_{\textit{slate}}=0$; $y_{\textit{government}}=2$; $y_{\textit{fiction}}=1$; $y_{\textit{travel}}=0$; $y_{\textit{facetoface}}=2$; $y_{\textit{nineeleven}}=0$; $y_{\textit{verbatim}}=1$.
In this way we generate spurious correlations within the dataset through sub-sampling to induce spurious correlations between $S$ and $Y$.
We show that the notion of $\rho_{\text{spurious}}$ in Equation 43 aligns with the notation in Definition 3.1.
Proposition J.1.
From the SCEs described above, we have that
$\mathbb{P}(Y=y_{s}\mid S=s)=\rho_{\text{spurious}}\;.$ (47)
for all spurious feature values $s$.
Proof.
This is clear considering Equation 44, where $\rho_{\text{spurious}}$ influences the chance that we sample $y_{s}$, i.e. the label spuriously correlated with feature value $s$. ∎
SCM for SMNLI. The DGP again induces a SCM that induces spurious correlations between spurious features $S$ and the label $Y$. The SCM has the same structure as in the SS dataset, and is given in Figure 16 where once again, $U$ again is some hidden confounding variable.
Figure 14: SMNLI SCM. SCM showing showing the spurious correlation present between the keyword feature $S$ and the label $Y$ of examples within the SMNLI dataset, induced through the described sub-sampling process of the MNLI dataset.
Independence conditions during training for FIT. As specified in Appendix C, we would like to have that $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ during training so that models trained via FIT can effectively learn to leverage focus instructions to make predictions based on specified features, where, again, $Y_{S}$ is the label spuriously associated to spurious feature value $S$. The results below give sufficient conditions for this to occur with respect to the DGP described in Figure 13.
Proposition J.2.
Assuming the SCEs given above and the corresponding DGP described in Figure 13, if we have a uniform label distribution $p(Y=y)$, and have that $\rho_{\text{spurious}}=1/3$ in the SMNLI training set, then we have that $Y\perp\!\!\!\perp S$.
Proof.
From Proposition J.1, we have that $\mathbb{P}(Y=y_{s}\mid S=s)=1/3$ when $\rho_{\text{spurious}}=1/3$ for any spurious feature value $s\in\text{Val}(S)$. Moreover, we have that from Equation 44
$\mathbb{P}(Y=y\mid S=s)$ $=\mathbb{P}(C=y\mid S=s)$ (48)
$=\underbrace{\mathbb{P}(U_{C}=1)}_{=\frac{2}{3}}\underbrace{\mathbb{P}(C=Y|S=s)}_{=\frac{1}{2}}$ (49)
$=\frac{1}{3},$ (50)
for each $y\in\text{Val}(Y)\setminus\{y_{s}\}$.
Using this alongside the assumed balanced label distribution, $\mathbb{P}(Y=y)=1/3$ for $y\in\text{Val}(Y)$, implies that we have $\mathbb{P}(Y=y|S=s)=\mathbb{P}(Y=y)$ for all $y\in\text{Val}(Y)$ and $s\in\text{Val}(S)$, which gives that $Y\perp\!\!\!\perp S$. ∎
Proposition J.3.
Assuming the SCEs given above and the corresponding DGP described in Figure 13, if we have a uniform label distribution $p(Y=y)$, and have that $\rho_{\text{spurious}}=1/3$ in the SMNLI training set, then we have that $Y_{S}\perp\!\!\!\perp C$.
Proof.
[…]
First, using that $\rho_{\text{spurious}}=1/3$, we have from Equation 44 that
$\mathbb{P}(C=c\mid S=s)$ =$=\begin{cases}1/3&\text{if }c=f(s),\\ 2/3\cdot 1/2&\text{else.}\end{cases}$ (51)
$=\frac{1}{3},$ (52)
[…]
Now we consider the marginal distribution of C. Let $|\text{Val}(S)|=k$ be the number of spurious feature values. Using the previous result, we have that:
$\mathbb{P}(C=c)$ $=\sum_{s}\underbrace{\mathbb{P}(C=c\mid S=s)}_{=1/3}\underbrace{\mathbb{P}(S=s)}_{1/k}$ (53)
$=\sum_{s}\frac{1}{3k}$ (54)
$=\frac{1}{3}.$ (55)
Therefore, we have that $\mathbb{P}(C=c\mid S=s)=\mathbb{P}(C=c)$, for all $c\in\text{Val}(C)$ and $s\in Val(S)$, which then implies that $S\perp\!\!\!\perp C$. Therefore, using that the spurious label $Y_{S}$ is a deterministic function of $S$, we have that $Y_{S}\perp\!\!\!\perp C$ under the assumptions within the proposition. ∎
Appendix K Spurious HANS dataset (SHANS)
We generate a binary NLI dataset, SHANS, with a known spurious feature. We do this considering the HANS dataset (McCoy, 2019). This is an NLI data set with two labels: entailment (0) and contradiction (1). This is an adversarial dataset designed to assess different NLI models’ reliance on spurious heuristics rather than on the underlying relationship between the premise and the hypothesis when making predictions. Specifically, the author’s consider three major categories of heuristics: lexical overlap heuristic (assuming that a premise entails from words within the hypothesis) , sub-sequence heuristic( assuming that the premise entails all any of its contiguous sub-sequences of words) and constituent heuristic (assuming that a premise entails a hypothesis that is any constituent within it’s syntactic parse tree).
Data-generating process (DGP). We consider a graphical model to describe the DGP of examples within the SHANS dataset. We use the following variables within our DGP:
  • $C$ NLI relationship between a premise and hypothesis pair, the core feature within this task, sampled from the original dataset.
  • $S_{\text{lex.}}$ spurious feature, here the presence of a hypothesis entirely made from words from the premise. This is a binary categorical variable (present/ not present).
  • $S_{\text{sub.}}$ spurious feature, here the presence of a hypothesis that is a contiguous subsequence of the premise. This is a binary category feature (present/ not present).
  • $S_{\text{const.}}$ spurious feature, here the presence of hypothesis that is a constituent/subtree of the premise. Here we have a binary variable (present/ not present).
  • $X$ example from the HANS dataset.
  • $Y$ final label for element $X$.
Figure 15: SHANS DGP. Graphical model showing the data generating process for modifying examples from the SHANS dataset to introduce a new spurious features $S_{\text{lex.}}$, $S_{\text{sub.}}$ and $S_{\text{const.}}$ which are encoded within the categorical spurious feature $S$ which represents one of these three heuristics.
The graphical model described by the DGP for producing the S-HANS dataset is given in Figure 15. Once again, this graphical model can be represented functionally as
$S$ $=f_{S}(U_{S});$ (56)
$C$ $=f_{C}(S,U_{C});$ (57)
$X$ $=f_{X}(C,E,U_{X});$ (58)
$Y$ $=f_{Y}(C,U_{Y}),$ (59)
where here we define $S$ to be a categorical feature over the set of the presence of each of the three heuristics introduced above which we denote, through overloaded notation, by $\mathcal{S}=\{s_{\text{lex.}},s_{\text{sub.}},s_{\text{const.}}\}$. More specifically, given the original dataset $\mathcal{D}$ that we are sub-sampling from, the functions that we use within the DGP for the S-HANS dataset are given by:
$S$ $\sim\texttt{Cat}(\mathcal{S}),$ (60)
$U_{C}$ $\sim\texttt{Ber}(\rho_{\text{spurious}});$ (61)
$C$$\sim\begin{cases}y_{S}&\text{ if }U_{C}=1;\\ \sim U(\text{Val}(Y)\setminus\{y_{S}\})&\text{ if }U_{C}=0;\end{cases};$ (62)
$X$ $\ \sim p_{\mathcal{D}}(\cdot|C,S)$ (63)
$Y$ $=C.$ (64)
Here, $\texttt{Cat}(\mathcal{S})$ is a uniform categorical distribution over $\mathcal{S}$ which effectively selects the precence of exactly one of the three spurious feature heuristics. We define $y_{S}$ to be the NLI label that a particular value of $S$ is spuriously correlated with by design. Moreover, $p_{\mathcal{D}}(x|C,S)$ is the conditional distribution over the dataset examples (premise-hypothesis pairs) that have NLI relationship $C$ and the presence of spurious heuristics $S$.
We consider the presence of each feature to be separate binary spurious features. The specific spurious correlations between heuristics and labels $Y$ are chosen to be: $y_{S_{\text{lex.}}=1}=0$; $y_{S_{\text{sub.}}=1}=0$; $y_{S_{\text{const.}}=1}=1$.
In this way we generate spurious correlations within the dataset through sub-sampling to induce spurious correlations between the heuristics and $Y$.
Transferred results from SMNLI. As we use the same DGP as for the SMNLI dataset described in Appendix K, all of the results that we have proven for SMNLI, translate to the SHANS dataset. In particular, we have that $\rho_{\text{spurious}}$ aligns with the notation in Definition 3.1, and that we have $Y\perp\!\!\!\perp S$ and $Y_{S}\perp\!\!\!\perp C$ under the assumptions of balanced label distributions and $\rho_{\text{spurious}}=1/2$ within the training set used for FIT training.
SCM for SHANS. The DGP again induces a SCM. In particular, considering $S$ as consiting of three binary spurious features $s_{\text{lex.}}$, $s_{\text{sub.}}$ and $s_{\text{const.}}$. The SCM has a similar structure to as in the SS and S-MNLI datasets, and is given in Figure 16 where once again, $U$ again is some hidden confounding variable.
Figure 16: SHANS SCM. SCM showing the spurious correlations present between the binary precence of heurstics features $S_{\text{lex.}}$, $S_{\text{sub.}}$ and $S_{\text{const.}}$ and the label $Y$ of examples within the S-HANS dataset, induced through the described sub-sampling process of the S-HANS dataset.
Appendix L FIT on SHANS
Here we give the results of performing SFT and FIT on the SHANS datasets.
Spurious HANS (SHANS) dataset. We generate binary NLI dataset sub-sampled from HANS (McCoy, 2019), a dataset designed to challenge NLI models by exposing common heuristics they rely on, such as lexical overlap (whether the hypothesis shares many words with the premise), sub-sequence (whether the hypothesis is a contiguous sub-sequence of the premise), and constituent (whether the hypothesis is a grammatical sub-structure of the premise). The presence of these heuristics are spuriously correlated with labels through sub-sampling of the presence of each of the heuristics from the original dataset. The degree of co-occurrence is governed by $\rho_{\text{spurious}}$, which varies according to the test sets described in Section 3. We ensure that $\rho_{\text{spurious}}$ is the same for all feature values within each dataset split. In particular, we set $\rho_{\text{spurious}}$ to be $0.5,0.5,0.9,0.25$ and $0.9$ on $\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{iid}}$, $\mathcal{D}_{\text{high}}$, $\mathcal{D}_{\text{low}}$ and $\mathcal{D}_{\text{flipped}}$ respectively.
Results. Figure 17 shows the focus accuracy results of performing SFT and FIT on the SHANS dataset for the Llama-3.1-8B-Instructmodel. As expected, the trained features show high focus accuracy. However, for non-trained features, we observe lower focus accuracy. This could be attributed to the overlapping nature of the heuristics in SHANS, which are often graded versions of each other with different levels of specificity. For instance, the sub-sequence heuristic can overlap with both lexical overlap and constituent heuristics (e.g., the example with Premise:“Before the actor slept, the senator ran” and Hypothesis: “The actor slept.” satisfies all three heuristics). This overlap likely confuses the model during generalisation, as it struggles to distinguish between heuristics not seen during training and those that are similar. These results suggest a potential limitation of FIT when dealing with features that are not sufficiently distinct or have significant overlap.
Figure 17: SHANS focus accuracies ($\uparrow$). Focus accuracy ($\mathcal{A}_{\text{focus}}$) of Llama-3.1-8B-Instructafter FIT on the SHANS dataset. Here, $C$ refers to the core feature (logical relationship between premise and hypothesis) and $S$ the spurious feature (heuristic used)